# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ 5.2 Local Environment and Retailers’ Incentives: Evidence from the Official Surveys
Next, I assess circumstances among the retailers using available survey reports on the credit institution. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^73] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those included retailers as well as wholesale and other small businesses. The number of respondents varied depending on the question items. The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930. In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). Thus, credit purchases in Tokyo were neither region- nor religion-specific customs.
@@
in the Retailers in Tokyo City
@@
stores for regular the Regular Customer (%) Cash Credit Monthly competitors
@@
Notes: This table summarizes the statistics on the sales in the $326$ retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. The number of stores reported the share of regular customers (column 2) is $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel categories, respectively. The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual sales. The definition of the social classes (“Upper”; “Middle”; “Other”) is not defined in the report. The share of sales method is based on the percentage of sales using each method in the annual sales. The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius. The department stores and retail markets are not included in this survey.
Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
@@
A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the retailers’ sales measured in this survey. [^74] This survey covered $939$ retailers, and its report was edited on December 31, 1935. I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels). The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses. I confirm that the main findings are materially similar. First, many annual sales were from their regular customers, most of whom were from the non-upper classes. $73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers. Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops. [^75] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used among greengrocers, $34$% of the sales were settled using credit. This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it. [^76] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. This confirms that the retailers were engaged in credit transactions with long-term relationships with their regular customers. [^77] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence for this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius. As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit. Online Appendix C.3 provides a set of regression-based evidence on the positive relationship between the use of credit purchases and the number of local competitors.
@@
Finally, I assessed the losses suffered by retailers for credit purchases. According to the 1930 Survey Report, $99$% of rice and $97$% of fuel retailers lost their sales. [^78] Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail. This implies that credit purchase was inherently a loss-making sales method for the retailers. However, these losses were usually marginal and capped at 5% of the total sales. For example, the available statistics suggest that $87$% of rice and $65$% of fuel retailers suffered losses amounting to $5$% of sales. While the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent. This seems to be a plausible sales strategy for retailers to manage adverse selection issues under imperfect consumer information.
@@
To summarize, the available documents show that retailers competed with other retailers in the same sales area. Under this circumstance, they provided regular consumers with the opportunity to trade on credit without interest for at least one month. The 1930 Survey Report indeed stated that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p. 415). This accessibility to credit purchases could help workers manage consumption after unexpected individual setbacks in the short term. [^79] Another potential explanation for the popularity of credit purchases might be theft. However, the number of thefts in $100$ people in 1919 was $0.5$ in Tsukishima, which is close to or even slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245). In fact, Tsukishima Police Department does not mention that Tsukshima was an area with a particularly high rate of theft (Tsukishima Police Department 1976, pp. 244–245). Therefore, the crime channel is not relevant to explain the use of credit purchases.
@@ 6 Conclusion
This study investigates the consumption-smoothing strategies used by working-class households in prewar Tokyo. I digitized the daily longitudinal budget survey on factory worker households conducted in Tsukishima in 1919. I assessed the potential measurement error due to the miss-assignments of paydays and proposed a way to avoid the assignment errors. The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks.
The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries. The elasticity for the rural households in Indian villages between 1975 and 1984 was reported to be approximately $0.5$. As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan. Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations. Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks. A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920.808080The estimates for India, Bangkok, and Osaka are obtained from Ravallion and Chaudhuri (1997), Townsend (1995), and Ogasawara (2024), respectively. My estimate from the adjusted monthly panels is slightly larger than this estimate. If taken literally, this implies that the Tsukishima factory workers were slightly more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work. This would add a new analytical perspective to the study of economic history.
Despite the vulnerability, the results from mechanism analyses show that credit purchases were commonly used among workers and played a role in mitigating adverse idiosyncratic shocks in several consumption subcategories, particularly food. My findings offer a new historical view of the interdependence between consumers and local retailers. As researchers build an understanding of local credit institutions we can comprehend regional consumption behavior and may understand how living standards improved in prewar Japan.
Appendices
@@ A.1 Development of the Cities
Notes: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo City (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan.
Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
Table A.1: City Characteristics:
Tokyo, Osaka, Kyoto, and Nagoya
Name of cities
Tokyo Osaka Kyoto Nagoya
Panel A: Population and Workers
Population 2,173,201 1,252,983 591,323 429,997
Male workers 768,633 476,229 189,922 138,275
Panel B: Industrial Structure (share in male workers)
Agriculture 1.0 0.8 2.2 1.8
Fisheries 0.1 0.1 0.0 0.4
Mining 0.4 0.3 0.2 0.2
Manufacturing 44.5 45.6 49.0 45.8
Commerce 32.4 34.0 32.3 29.6
Transport 7.6 10.6 5.5 8.0
Public service and professions 11.3 6.7 9.0 11.3
Housework 0.3 1.8 0.1 0.0
Other industry 2.4 0.1 1.5 2.8
Panel C: Local Infrastructures
Number of electrical lights per people in 1921 0.9 0.8 0.9 0.9
Accessibility to modern water supply (%) 78.9 80.5 46.5 39.4
Panel D: Meteorological Features
January 1918 – December 1923
Average daily temperature 14.1 15.4 14.1 14.6
Average daily precipitation 4.7 3.9 4.6 4.8
February – April only
Average daily temperature 8.0 8.8 7.8 8.3
Average daily precipitation 4.2 3.4 4.1 4.3
Notes:
Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed.
Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation.
Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census. The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%).
Panel D: The average daily temperature and precipitation between January 1, 1918, and December 31, 1923, and between February 1, 1920, and April 30, 1920 are listed.
Sources: The number of people in 1920 is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of male workers and their industrial structures are from the Statistics Bureau of the Cabinet (1929a, pp. 20–21; 26–29). The number of houses accessible to the water supply is from the Water Works Association (1922, pp. 69–70). The total number of households is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of electrical lights is from the Ministry of Posts and Telecommunications (1922, p. 558) The daily temperature and precipitation are from the Japan Meteorological Agency (database).
Figure A.1a illustrates the developments of the large cities in Japan between 1889 and 1920. Table A.1 summarizes the scales, industrial structures, infrastructures, and meteorological features of the large cities.
@@ A.2 Financial Institutions in the Cities
Several financial institutions were available for working-class households in the cities circa 1920.
A representative savings institution was the postal savings (yūbin chokin). Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was approximately $10$% in the 1910s (Tanaka 2018, p. 29). However, the number of postal savings accounts per total population in the early 1910s exceeded $20$% (Tanaka 2018, p. 33). James and Suto (2011) indeed show that the saving rate in prewar Japan was higher than that of the US. As shown in Panel A of Table A.2, approximately $1.4$ million people had accounts in Tokyo in 1920. This covers approximately $67$% of the entire population in Tokyo City and the average savings amount per capita was $64.5$ yen. In Kyobashi Ward, the average figure for the postal savings accounts was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp. 892–893). This amount was less than the average monthly earnings of the skilled-factory workers (Section 3.2). Unfortunately, the statistics on postal savings by depositors’ occupation are unavailable. [^81] There are some statistics for the whole of Japan. According to the Postal Savings Bureau (1924, p. 72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount was $5.5$%). However, this implies that, while a large proportion of the workers might have had accounts, postal savings may not have provided sufficient temporary income.
Savings bank (chochiku ginkō) was another popular savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81). Panel A of Table A.2 shows that the number of people who had savings accounts was approximately $1.46$ million, accounting for approximately $67.3$% of the total people in Tokyo City at that time. Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors. Table A.3 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B). In Kyobashi, the commerce and miscellaneous sectors shared approximately $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%. This is unchanged if I use the amounts of deposits to calculate the shares. The Tokyo Institute for Municipal Research confirmed that savings banks were designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p. 81). Thus, skilled factory worker households—like the THBS households—were less likely to use the savings banks. This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices. [^82] There were 123 savings banks in Tokyo City in 1922, and ten of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p. 76). However, Tsukishima—a representative manufacturing area—did not have savings banks at that time.
Mutual loan association (mujin) was another type of savings institution. The members of this association paid a fixed amount of money several times, and each member withdrew the total amount of payments each time under specific rules. The lotteries or bidding system decided which members could withdraw money each time. This means that the workers could not use this institution to manage acute adverse income shocks. [^83] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin. There was a different type called sueoki chokin (or tsumitate kai) from which people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp. 3–4). However, these were also not flexible for compensating for income shocks. Instead, these were used by small business owners and merchants to generate money. A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo City (Tokyo City Office 1935, p. 25). Moreover, the Tokyo Chamber of Commerce stated that the mutual loan association was used for business and rarely for household consumption (Tokyo Chamber of Commerce 1918, p. 121).
Cooperative societies were also available for worker households. Cooperative societies in Tokyo used credit as their payment method—households could buy goods in advance and pay at the end of the month or twice (in the middle and at the end of) a month. The commodity prices were similar to or slightly lower than those in the retailers (Central Federation of Industrial Associations 1925, pp. 18–19). However, the cooperatives covered a small proportion of citizens. There were $26$ cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). Among these $23$ cooperatives, two were large societies with approximately $4,000$ members, and the others were small societies with less than $100$ members. Given that the number of male workers was more than $768$ thousand (Table A.1), these account for only a few percentage points of the total workers in the city. In fact, the THBS households rarely used the cooperatives.
An alternative borrowing institution was credit purchase in retailers. The details of the credit purchase are little documented, and the systematic statistics on credit purchases are also unavailable. I surmise that this was because the information on the transactions was personal. Despite this, a few available documents show that credit purchases were used in various retailers at that time (Section 5). The THBS households frequently used credit for purchases in various retailers for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines.
Pawnshops were the most popular lending institution among low-income working-class households. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles pawned were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895. Thus, the average redemption rate was substantially high under the lower interest rates. Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp. 888–889). The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures were more than three times greater than the entire population in Tokyo City. This means that pawnshops were more accessible to workers than other institutions. As described in the main text, there were $13$ pawnshops in Tsukishima, whereas there were five moneylenders.
However, the inexpensive articles pawned meant that pawnshops were available only for short-run necessities. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was approximately $10$% of the monthly earnings of the THBS households. While the pawnshop was a popular lending institution, it could not offer enough temporary income for relatively large idiosyncratic shocks. According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of users in the agriculture, commerce, and manufacturing sectors was $1.7$%, $19.5$%, and $15.6$%, respectively. In turn, the most representative users were the day laborers and workers classified as “miscellaneous,” accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p. 25). In fact, the Tokyo City Office stated that the pawnshops were mainly used among low-income (poor) working-class households (Tokyo City Social Affairs Bureau 1921, p. 9). This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
Another lending institution was the money lenders (kinsen kashitsuke gyō). Generally, business owners used money lenders. The interest rates of the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp. 184; 248). Tokyo Institute for Municipal Research was indeed concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp. 106–108). Given this fact, while there were two money lenders in Tsukishima, they were less likely to be used by the factory worker households.
Credit union (shinyō kumiai) was another institution available for working-class households. In Tokyo City, there were $215$ unions in 1923. According to the survey of the $184$ unions among them, however, the number of union members in the manufacturing sector was only $5,662$ (Tokyo Institute for Municipal Research 1925c, p. 20). This accounts for only a few percentage points of the male workers in the manufacturing sector in Tokyo measured in the population census of 1920 (Online Appendix Table A.1).
Table A.2: Financial Institutions in Tokyo, Osaka, Kyoto, and Nagoya Cities in 1920
@@
Notes:
1. Panel A: The statistics on postal savings were measured at the end of March 1921. The statistics on savings banks were measured at the end of 1920.
[…]
@@
Pane A: Entire Tokyo City
@@
Pane B: Kyobahi Ward
@@
Notes: This table summarizes the number of depositors and the total deposits in the savings banks by occupation. In the document, the miscellaneous category is not clearly defined but named as “Miscellaneous (zatsu gyō).” These statistics were measured on December 31, 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively.
Source: Tokyo City Office (1922a, pp. 864–865).

### v14

@@ 5.2 Local Environment and Retailers’ Incentives: Evidence from the Official Surveys
Next, I examine the circumstances among retailers using a survey report published by the Tokyo City Office in 1930 (hereafter, the 1930 Survey Report). [^58] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1), including retailers, wholesalers, and other small businesses. The number of respondents varies by question. The statistical table on payment methods includes $2,654$ respondents, of whom $1,197$ were retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of retailers in Tokyo circa 1930. Although a similar survey around 1920 is unavailable, the industrial and household structures in 1920 and 1930 were generally comparable (Online Appendix Table C.6). Thus, the 1930 report can serve as a reference to illustrate how credit purchases functioned in the city. According to the report, credit purchases were a popular payment method across consumer transactions: 9999%($155/157$) of grain retailers, 8181%($35/43$) of fish retailers, 7878%($64/82$) of greengrocers, and 9494%($45/48$) of fuel retailers used credit purchases as part of their selling methods (Tokyo City Office 1932b, pp. 260; 263). This widespread use indicates that credit purchases were neither region- nor religion-specific customs.
@@
among Retailers in Tokyo City
@@
stores for regular regular customers (%) Cash Credit Monthly competitors
customers (%) Upper Middle Other purchase installment
@@
Notes: This table summarizes sales statistics for $326$ retailers in Tokyo surveyed in 1935, covering white rice, fish, fruits and vegetables, and firewood and charcoal. The number of stores reporting the share of regular customers (Column 2) is $90$, $56$, $61$, and $93$ for the white rice, fish, fruits and vegetables, and fuel categories, respectively. The “share of sales for regular customers” refers to the percentage of annual sales made to regular customers. Social classes (“Upper,” “Middle,” and “Other”) are not defined in the report. The share of sales by method is based on the percentage of annual sales conducted using each method. The number of competitors indicates peers within a three-cho radius (approximately $327$ meters). Department stores and retail markets are excluded from this survey. Source: Tokyo Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
@@
A survey of retailers in 1935, published by the Tokyo Chamber of Commerce and Industry, provides detailed information on sales practices. Table 9 summarizes these statistics. [^59] The survey covered $939$ retailers, and the report was finalized on December 31, 1935. It includes detailed questions regarding sales and purchases (Tokyo Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.3 decomposes the statistics in Table 9 by business scale, confirming that the main findings are materially similar. First, a large share of annual sales come from regular customers, most of whom belonged to non-upper classes. 7373% of annual sales by rice retailers come from regular customers, and more than 8080% of these customers were from non-upper classes, including factory workers. Second, credit purchase was a popular sales method, particularly among rice and fuel retailers: more than 7070% of sales in both categories were conducted on credit. [^60] These figures are systematically lower than those in the 1930 Survey Report. Both surveys, however, show the same general tendency toward credit purchases; differences may reflect variations in the samples, retailer category definitions, or the definition of credit purchase itself. Cash was also accepted by greengrocers, but 3434% of their sales were on credit. This confirms that retailers engaged in credit transactions, often maintaining long-term relationships with regular customers. [^61] A 1924 textbook for retailers explains the psychology behind these long-term relationships: credit purchases are convenient for customers from a budget perspective and confer “vanity” or status as regular customers (Shimizu 1924, pp. 115–116). Retailers also commonly used credit purchases to buy source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157), making credit a familiar payment method for both retailers and consumers. The final column of Table 9 provides suggestive evidence on competition. The survey asks whether the sales area is competitive and, if so, how many competitors are located within a three-cho (approximately $327$ meters) radius. Rice and fuel stores have two to three more competitors than retailers selling fish or fruits and vegetables. Given the relatively high use of credit among rice and fuel retailers, this suggests that they may have responded to competition by extending credit to their customers. Online Appendix C.5 presents regression-based evidence of a positive relationship between credit use and the number of local competitors.
@@
Finally, I assess the losses suffered by retailers from unpaid credit purchases. According to the 1930 Survey Report, 9999% of rice retailers and 9797% of fuel retailers experienced losses on credit sales. [^62] Online Appendix C.6 summarizes detailed statistics on losses from credit purchases in the 1930 Survey Report. This indicates that credit purchases were inherently loss-making for retailers. However, these losses were generally small, typically capped at 5% of total sales. For example, available statistics show that 8787% of rice retailers and 6565% of fuel retailers suffered losses amounting to 55% of sales. Although retailers could not avoid losses from credit sales, they may have set prices to anticipate these expected losses. This appears to be a plausible strategy for managing adverse selection under imperfect consumer information (Section 5.1).
@@
In summary, the available documents show that retailers competed with others in the same sales area and provided regular customers the opportunity to trade on credit without interest for at least one month. The 1930 Survey Report noted that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p. 415). This informal credit system could help workers manage consumption following unexpected individual setbacks in the short term. [^63] Another potential explanation for the popularity of credit purchases might be theft. However, the number of thefts per $100$ people in 1919 was $0.5$ in Tsukishima, close to or slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245). The Tsukishima Police Department does not report that the area had an unusually high rate of theft (Tsukishima Police Department 1976, pp. 244–245). Therefore, the crime channel does not explain the use of credit purchases.
@@ 6 Conclusion
This study investigates the consumption-smoothing strategies of skilled factory-worker households by digitizing the longitudinal budget survey conducted in Tsukishima in 1919. Despite this narrow geographical focus, the analytical sample represents average skilled factory-worker households in this representative manufacturing area of Tokyo.
The estimated elasticity for total consumption indicates that households experienced measurable consumption losses when facing idiosyncratic income shocks. The estimated elasticity for total expenditure ($0.43$) is similar to estimates found in developing countries. The elasticity for rural households in Indian villages between 1975 and 1984 is reported to be approximately $0.5$. In Bangkok, which had a GDP per capita similar to that of Japan in 1920, the estimated elasticity for urban households is $0.4$ between 1975 and 1990.646464The estimates for India, Bangkok, and Osaka are from Ravallion and Chaudhuri (1997), Townsend (1995), and Ogasawara (2024), respectively.
A recent study reports an income elasticity of $0.39$ for factory-worker households in Osaka around 1920. My estimate from the monthly panels is slightly higher. Taken literally, this suggests that Tsukishima factory workers were slightly more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to greater availability of risk-coping institutions. A limitation of this study is that it does not conduct inter-city and cross-national comparisons of income elasticities, nor does it analyze the institutional characteristics affecting consumption; these are left for future research.
Despite this vulnerability, results from the mechanism analyses show that credit purchases were commonly used by workers and helped mitigate adverse idiosyncratic shocks in several consumption subcategories, particularly food. These findings offer a new historical perspective on the interdependence between workers and local retailers. As researchers develop a better understanding of local credit institutions, we may gain insight into regional consumption behavior and the factors that improved living standards in prewar Japan.
@@ A.1 Development of the Cities
Notes: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo City (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan. Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
Figure A.1a shows the developments of the large cities in Japan between 1889 and 1920.
@@ A.2 Financial Institutions in the Cities
Table A.1: Financial Institutions in Tokyo, Osaka,
Kyoto, and Nagoya Cities in 1920
@@
Notes: Panel A: The statistics on postal savings were measured at the end of March 1921. The statistics on savings banks were measured at the end of 1920. Panel B: Pawnshops in Tokyo, Osaka, and Kyoto were measured in 1920. Those for Nagoya (suggested in $\dagger$}) were measured in 1923 because the statistics in the former years are unavailable.
@@
A representative savings institution was postal savings (yūbin chokin). Panel A of Table A.1 shows that approximately $1.46$ million people held postal savings accounts in Tokyo in 1920, covering about 6767% of the city’s population. However, the average savings per capita was insufficient. In Kyobashi Ward, the average balance of postal savings accounts was only $44.5$ yen in 1919 (Tokyo City Office 1921, pp. 892–893). This amount is lower than the average monthly earnings of skilled factory workers (Section 3.2), suggesting that postal savings may not have provided adequate temporary income. [^65] Although statistics on postal savings by depositors’ occupation are unavailable, some nationwide data exist. According to the Postal Savings Bureau (1924, p. 72), depositors in the manufacturing sector accounted for 4.84.8% of the total, and their share based on deposit amounts was 5.55.5%).
Savings banks (chochiku ginkō) were another common savings institution (Ito and Saito 2019, pp 79–81). Panel A of Table A.1 shows that the number of people with savings accounts was approximately $1.46$ million, representing approximately 6767% of Tokyo’s total population. Although this figure is similar to that for postal savings, savings banks were designed as financial institutions for low-income workers (Tokyo Institute for Municipal Research 1925b, p. 81). [^66] There were 123 savings banks in Tokyo City in 1922, and 10 of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p. 76). However, Tsukishima—a representative manufacturing area—did not have savings banks at the time. Table A.2 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B). In Kyobashi, the commerce and miscellaneous sectors together accounted for about 84.884.8% of all depositors, while the manufacturing sector accounted for only 11.511.5%. These shares remain unchanged when calculated based on deposit amounts.
Mutual loan associations (mujin) represent another type of savings institution. Members contribute a fixed amount of money over several periods, and, under specific rules, each member withdraws the total contributions once during the cycle. A lottery or bidding system determines withdrawal order, which means that workers could not rely on these associations to manage sudden adverse income shocks. [^67] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin. There was a different type called sueoki chokin (or tsumitate kai) from which people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp. 3–4). However, these were also not flexible for compensating for income shocks. Instead, they were mainly used by small business owners and merchants to generate revenue (Tokyo Chamber of Commerce 1918, p. 121). Workers in the manufacturing sector accounted for only 1010% of all mutual loan associations in Tokyo City (Tokyo City Office 1935, p. 25).
Cooperative societies were also available for workers. In Tokyo, these societies used credit as their payment method; households could purchase goods in advance and settle their payments at the end of the month or in two installments, mid-month and month-end. Commodity prices in these cooperatives were similar to, or slightly lower than, those in retail shops (Central Federation of Industrial Associations 1925, pp. 18–19). In 1924, Tokyo Prefecture had $26$ cooperative societies, of which 23 were organized by citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). Among these $23$ cooperatives, two were large, with approximately $4,000$ members each, while the rest were small, with fewer than $100$ members. Given that the number of male workers exceeded $768$ thousand (Statistics Bureau of the Cabinet 1929a, pp. 20–21; 26–29), these cooperatives accounted for only a small share of all workers in the city.
Another borrowing option was purchasing on credit from retailers. Details on credit purchases are poorly documented, and systematic statistics are unavailable because transaction records are personal. Nevertheless, a few existing sources indicate that credit purchases were widely used across various retailers (Section 5). The THBS households frequently used credit to buy goods, including rice, fish, vegetables, alcohol, firewood, charcoal, and medicine, from different retailers.
@@
Panel A: Entire Tokyo City
@@
Panel B: Kyobashi Ward
@@
Notes: This table summarizes the number of depositors and the total deposits in the savings banks by occupation. In the document, the miscellaneous category is not clearly defined but named as “Miscellaneous (zatsu gyō).” These statistics were measured on December 31, 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively. Source: Tokyo City Office (1922a, pp. 864–865).
According to the Tokyo City Office, pawnshops were primarily used by low-income workers (Tokyo City Social Affairs Bureau 1921, p. 9). Lenders did not need to assess borrowers’ creditworthiness, and borrowers faced limited risk of heavy debt because most pawned items were inexpensive clothes (Shibuya et al. 1982). The average redemption rate was high, supported by relatively low interest rates regulated under the Pawnbroker Regulation Act of 1895. Among the $88$ pawnshops in Kyobashi, the average redemption rate was approximately 9494% (401,222401,222/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp. 888–889). In Tokyo, overall, the number of pawnshops was $1,334$ in 1918 and $1,261$ in 1919, with total transactions of $8,226,883$ and $7,573,406$, respectively. These figures exceed three times the total population of Tokyo City, indicating that pawnshops were more accessible to workers than other financial institutions. Figure A.2 shows that Tsukishima had $13$ pawnshops compared to five moneylenders.
However, because the items pawned were inexpensive, pawnshops primarily served short-term needs. For example, in Kyobashi, the average loan per transaction was $7.8$ yen, approximately 1010% of the monthly earnings of the THBS households. Statistics on pawnshop users in Tokyo Prefecture in 1923 show that 1.71.7% worked in agriculture, 19.519.5% in commerce, and 15.615.6% in manufacturing. The largest group of users consisted of day laborers and workers classified as “miscellaneous,” accounting for 63.463.4% (Tokyo Institute for Municipal Research 1926, p. 25).